OpenStudy (gabylovesyou):

solve the proportion 2x/5x + 6 = 2/x is the answer x = -1 and x = 6 @ParthKohli @hartnn @ganeshie8

Parth (parthkohli):

Yes

ganeshie8 (ganeshie8):

how do u get all correct answers always ! awesome :)

OpenStudy (gabylovesyou):

yaaay and sorry if i bother i just wanna check my answers :P

Parth (parthkohli):

BTW, you can check your answers through http://wolframalpha.com

OpenStudy (gabylovesyou):

ok thanks :)

hartnn (hartnn):

why do i get x=5/16 ?

OpenStudy (gabylovesyou):

i have no clue lol

Parth (parthkohli):

@hartnn because you're solving the wrong equation

Parth (parthkohli):

You're solving\[\dfrac{2x}{5x} + 6 = \dfrac{2}{x}\]

hartnn (hartnn):

yes^

Parth (parthkohli):

lol, you've been using Wolfram all the time?!

Parth (parthkohli):

@hartnn But the question is\[\dfrac{2x }{5x + 6} = \dfrac{2}{x}\]

hartnn (hartnn):

no...for this one to verify... okk....missing paranthesis creates confusion....